### v1

@@ 4.5 Influence of Model Size
ViT-B 58.5 55.6 61.4
ViT-L 62.3 59.0 65.6
ViT-H 63.2 60.0 66.5
@@
Figure 4: Visualization of RefSAM on Ref-DAVIS17. The text with a specific color corresponds to the object mask of the same color.
@@
Figure 5: Visualization of different models on Ref-DAVIS17. From left to right: RefSAM, ReferFormer [4], SAM-Track [21] + Ground DINO [25], and PerSAM [50] + Ground DINO. The text with a specific color corresponds to the object mask of the same color.
4.6 Visualization Results
We show the visualization results of our RefSAM model in Figure 4. It can be seen that RefSAM is capable of effectively segmenting and tracking the referred object even in challenging scenarios, such as variations in person poses, and occlusions between instances. Furthermore, we present the results of differnt models in Figure 5. It is clear that our RefSAM demonstrates significantly enhanced cross-modal understanding capability, particularly evident in handling vague language descriptions (Figure 4(a)) and resolving appearance ambiguity among similar objects (Figure 4(c)).
Table 5: Number of learnable parameters of different Models.
@@ 4.6 Visualization Results
RefSAM ViT-B 4.8
RefSAM ViT-L 4.8
RefSAM ViT-H 4.8
Table 6: Inference speed of different models.
@@
RefSAM ViT-H 4.6
RefSAM ViT-L 7.0
RefSAM ViT-B 9.2
4.7 Model Complexity Analysis
In this section, we perform an analysis of the complexity of various models. As depicted in Table 5, our model exhibits a notably smaller number of learnable parameters in comparison to all other models, highlighting the inherent nature of parameter-efficient tuning employed during the training of our RefSAM model. Furthermore, Table 6 presents the inference speed of various models. It is apparent that our RefSAM exhibits a marginally slower runtime compared to ReferFormer, yet outperforms SAM-Track and PerSAM in terms of inference speed. When employing a ViT-B as the visual encoder, RefSAM achieves a faster inference speed of 9.2 FPS, outperforming all other methods.
5 Conclusion
This study pioneers the adaptation of the foundational segmentation model, Segment Anything Model, to the referring video object segmentation task. We propose the novel RefSAM model, which incorporates efficient designs to effectively bridge the semantic gap between the visual and language domains, significantly enhancing the cross-modal understanding capability of SAM. By employing a parameter-efficient tuning strategy, RefSAM achieves efficient training through adjustments to a small number of learnable parameters. Extensive experiments on Ref-Youtu-VOS and Ref-DAVIS17 datasets validate the superior effectiveness of our RefSAM model compared to existing methods.
[…]

### v3

@@ IV-D3 Impact of Different Text Encoders
T5-3b [50] (Tab. III) 60.3 57.3 63.3
Roberta-base [61] 58.7 55.6 61.7
Clip-large [62] 58.3 56.1 61.3
IV-D4 Influence of Adapter
In this section, we examine the impact of the Adapter. All experiments use the data augmentation described in Sec. IV-B2 and the model employs Cross-modal MLP and HDA. The results on Refer-Youtube-VOS and Ref-DAVIS17 are shown in VI. We conducted experiments on ViT-B. As shown, using ITM or the Adapter improves performance on both Refer-Youtube-VOS and Ref-DAVIS17. Moreover, using both ITM and the Adapter together results in even better performance. Using ViT-B on Refer-Youtube-VOS as an example, not using ITM or the Adapter yields a J&F-Mean score of 52.2 points. Adding ITM or the Adapter increases the score by 2.5 or 4.0 points, respectively. Finally, using both ITM and the Adapter together achieves a J&F-Mean score of 58.4 points, an increase of 6.2 points.
TABLE VI: Ablation study of adapter.
ITM Adapter Refer-Youtube-VOS Ref-DAVIS17
[…]
52.2 52.0 52.5 60.3 57.3 63.3
54.7 53.9 55.5 62.5 59.5 65.4
56.2 55.3 57.1 61.9 58.5 65.5
58.4 57.4 59.4 62.1 59.0 65.3
@@ IV-D4 Influence of Adapter
Figure 6: Visualization of RefSAM’s results on Ref-DAVIS17. The text with a specific color corresponds to the object mask of the same color.
(a) the glasses worn by a man standing in front of a red display car.
(b) a small fox like dog on the right side of a sofa.
(c) the black truck with red/white and blue is moving down the road to the right with a crowd behind it.
Figure 7: Visualization of the impact of HDA in our RefSAM on Refer-Youtube-VOS. The first and second columns show the cases without using HDA, and the third and fourth columns are the showcases using HDA.
@@
Figure 8: Visualization of the results of different models on Ref-DAVIS17. From left to right: our RefSAM, ReferFormer [4], SAM-Track [17] + Ground DINO [21], and PerSAM [45] + Ground DINO. The text with a specific color corresponds to the object mask of the same color.
IV-D5 Influence of Model Size
We then investigate the influence of model size, specifically the performance of using different sizes of Visual Encoder, including ViT-B, ViT-L, and ViT-H. All the models employ the Cross-modal MLP, HDA, Adapter, and ITM. The data augmentation described in Sec. IV-B2 is also used.The results on Refer-Youtube-VOS and Ref-DAVIS17 are shown in Table VII. The findings indicate that employing a larger backbone for visual feature extraction yields superior performance compared to a smaller one, attributed to the enhanced representation capacity of larger vision transformers. Furthermore, the results exhibit a consistent improvement in performance with increasing model size, highlighting the good scalability of our RefSAM model.
TABLE VII: Ablation study of the Visual Encoder.
Visual Encoder Refer-Youtube-VOS Ref-DAVIS17
[…]
ViT-B 58.4 57.4 59.4 62.1 59.0 65.3
ViT-L 62.2 60.8 63.6 70.5 66.7 74.3
ViT-H 67.6 65.8 69.4 71.9 68.4 75.5
IV-E Visualization Results
We show the visualization results of our RefSAM model in Figure 6. It can be seen that RefSAM is capable of effectively segmenting and tracking the referred object even in challenging scenarios, such as variations in person poses, and occlusions between instances. Besides, we also visualize the impact of using HDA in Figure 7. It can be observed that leveraging HDA greatly enhances the model’s capabilities in handling tiny objects and occlusions and providing more accurate segmentation masks, especially when dealing with complex text prompts. Furthermore, we present the results of different models in Figure 8. It is clear that our RefSAM demonstrates significantly enhanced cross-modal understanding capability, particularly evident in handling vague language descriptions (Figure 8(a)) and resolving appearance ambiguity among similar objects (Figure 8(c)).
TABLE VIII: Results on three representative RIS datasets.
Methods Visual Backbone Textual Encoder RefCOCO RefCOCO+ G-Ref
Val Test A Test B Val Test A Test B Val Test
oIoU
CGAN [63] DeepLab-R101 Bi-GRU 64.86 68.04 62.07 51.03 55.51 44.06 51.01 51.69
ISFP [64] Darknet53 Bi-GRU 65.19 68.45 62.73 52.70 56.77 46.39 52.67 53.00
LTS [65] Darknet53 Bi-GRU 65.43 67.76 63.08 54.21 58.32 48.02 54.40 54.25
ReSTR [66] ViT-B TX 67.22 69.30 64.45 55.78 60.44 48.27 - -
LAVT [67] Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
RefSegformer [68] Swin-B BERT 73.22 75.64 70.09 63.50 68.69 55.44 62.56 63.07
RefSAM ViT-H T5-3b 74.89 77.49 70.90 65.88 71.60 57.26 66.46 67.23
mIoU
VLT [40] Darknet53 Bi-GRU 65.65 68.29 62.73 55.50 59.20 49.36 52.99 56.65
CRIS [69] CLIP-L CLIP-L 70.47 73.18 66.10 62.27 68.06 53.68 59.87 60.36
SeqTR [70] Darknet53 Bi-GRU 71.70 73.31 69.82 63.04 66.73 58.97 64.69 65.74
RefTR [71] ResNet101 BERT 74.34 76.77 70.87 66.75 70.58 59.40 66.63 67.39
LAVT [67] Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
RefSAM ViT-H T5-3b 78.29 80.65 75.07 71.58 76.03 64.35 71.45 71.98
IV-F Model Complexity Analysis
In this section, we perform an analysis of the complexity of various models. As depicted in Table IX, our model exhibits a notably smaller number of learnable parameters in comparison to all other models, highlighting the inherent nature of parameter-efficient tuning employed during the training of our RefSAM model. Furthermore, Table X presents the inference speed of various models. It is apparent that our RefSAM exhibits a marginally slower runtime compared to ReferFormer and SgMg, yet outperforms SAM-Track and PerSAM in terms of inference speed. When employing a ViT-B as the visual encoder, RefSAM achieves an inference speed of 9.5 FPS.
TABLE IX: The number of learnable parameters of different models.
@@ IV-F Model Complexity Analysis
RefSAM ViT-B 9.9
RefSAM ViT-L 19.1
RefSAM ViT-H 33.0
TABLE X: Inference speed of different models.
@@
RefSAM ViT-H 4.8
RefSAM ViT-L 6.5
RefSAM ViT-B 9.5
IV-G Results on RIS Datasets
Our model can also be utilized for RIS. We evaluate our model on three representative datasets: RefCOCO [72], RefCOCO+ [72] and G-Ref [73, 74]. RefCOCO comprises 142K referring language expressions describing 50K objects in approximately 20K images, while RefCOCO+ consists of 141K referring language expressions for 50K objects in roughly 20K images. G-Ref includes 104K referring language expressions for around 55K objects in about 27K images. Following  [67], we train our model for 40 epochs using a batch size of 16 for each dataset. We use mean Intersection-over-Union (mIoU) and overall Intersection-over-Union (oIoU) as the evaluation metrics. For all experiments on the RIS datasets, we employ Cross-modal MLP, HDA and Adapter, and use the data augmentation described in Sec. IV-B2. The results are shown in Table VIII. The proposed RefSAM achieves competitive results in terms of oIoU and mIoU. This highlights the versatility of our RefSAM and its exceptional performance for RIS.
V Conclusion
This study pioneers the adaptation of the foundational segmentation model, Segment Anything Model, to the referring video object segmentation task. We propose the novel RefSAM model, which incorporates efficient designs to effectively bridge the semantic gap between the visual and language domains, significantly enhancing the cross-modal understanding capability of SAM. By employing a parameter-efficient tuning strategy, RefSAM achieves efficient training through adjustments to a small number of learnable parameters. Extensive experiments on two representative referring video object segmentation datasets including Refer-Youtube-VOS and Ref-DAVIS17 datasets, as well as three referring image segmentation datasets validate the superior effectiveness of our RefSAM compared to existing methods.
VI Data availability statements
Refer-Youtube-VOS is available for download from The Large-scale Video Object Segmentation Challenge  [^3] https://youtube-vos.org/ at https://doi.org/10.1007/978-3-030-58555-6_13[27]. Ref-DAVIS17 is available for download from International Max Planck Research School on Trustworthy Computing (IMPRS-TRUST)  [^4] https://www.mpi-inf.mpg.de/departments/computer-vision-and-machine-learning/research/video-segmentation/video-object-segmentation-with-language-referring-expressions at https://doi.org/10.1007/978-3-030-20870-7_8[25]. RefCOCO and RefCOCO+ are available for download from UNC Vision Recognition Group at https://doi.org/10.1007/978-3-319-46475-6_5[72]. G-Ref is available for download from UNC Vision Recognition Group at https://doi.org/10.1109/cvpr.2016.9[74].